OpenStudy (anonymous):

Bryce insists that (2+x)/8 is equivalent to (1+x)/4. What mistakes do you think is being made and how could you demonstrate to Bryce that the two expressions are not equivalent?

myininaya (myininaya):

(1+x)/4 *2/2=2(1+x)/4

myininaya (myininaya):

oops that should be 8 on the bottm

myininaya (myininaya):

(1+x)/4 *2/2=2(1+x)/8 2(1+x) is not the same as 2+x

OpenStudy (anonymous):

\[\frac{2+x}{8}=\frac{1+x}{4}\rightarrow \frac{8+4x}{8}=1+x \rightarrow 8+4x=8+8x \] \[4x=8x \rightarrow x=2x\] Which doesn't make sense since this says the variable x is equal to twice its value

myininaya (myininaya):

they are only equal when x=0

OpenStudy (anonymous):

That's true

myininaya (myininaya):

you can also graph these two and show that they aren't the same line

OpenStudy (anonymous):

you can do that as well

myininaya (myininaya):

you can also show him a counterexample to his logic

myininaya (myininaya):

by saying if we let x be 1 here what happens we ger 3/8 but over we get 2/4 2/4 does not equal to 3/8 sorry bryce

OpenStudy (anonymous):

The math i did earlier covers all examples for any value of x, Which demonstrate that the two expressions are not equal when x is greater than or less than 0, it is only true when x=0 as myininaya stated earlier